What is the probability of selecting two cards of different suits with replacement? What about without replacement? What if we want to select two cards of different ranks instead?
Answer #1

Ans:

With replacement:

probability of selecting two cards of different suits=(13/52)*(39/52)=0.1875

Without replacement:

probability of selecting two cards of different suits=(13/52)*(39/51)=0.1912

Now cards with different ranks:

probability of selecting two cards of different ranks with replacement=(4/52)*(48/52)=0.0710

probability of selecting two cards of different ranks without replacement=(4/52)*(48/51)=0.0724